OpenStudy (anonymous):

match 2/3 1/9 60% 0.35 1/80 4/25 110% 125% 0.08 0.125 with 2/25 1 1/4 1.25% 16% 1.1 7/20 66.6 (with a line over the last 6) 3/5 1/8 0.1 (with a line over the 1) please and thank you! :)

OpenStudy (anonymous):

@agent0smith @Ashleyisakitty @allie_bear22 @Ammarah @bahrom7893 @blurbendy @bookworm00981 @Brad1996 @charlotte123 @CalebBeavers @caitlinnr14 @cantthink96 @Directrix @DemolisionWolf @demonguy @DeadShot @Easyaspi314 @Euler271 @emilyhaddad @Emily778 @Firejay5 @francoisderozay @furything @frankhecker @Gatorgirl @GABI_J @gripsocks @Hero @helpme1.2 @halorazer @helppls123 would any of you guys know? :)

OpenStudy (anonymous):

i dont sorry

OpenStudy (anonymous):

play with your calculator to try and narrow some down.

OpenStudy (anonymous):

wow, sry. I don't know these.

OpenStudy (anonymous):

@IanK16. @INewton @ittsninaa @imawesome101 @jigglypuff314 @jaredfleagle @juciystar234 @jpsmarinho @katlin95 @KattGirl14 @Kristen17 @kaylalynn @Loser66 @lowcard2 @LaynaMae @Little_Kay @myininaya @msumner @marylou004 @mangorox @Nuke_ur_friend @nishia @Nutcase482 @Niniguma @OrionsBelt @Objection @opthompson @poopsiedoodle @Piggiepigpigs @plohr @PurplePanda312 @ranga @river_song @Rose96<3 @RainbowsxD @satellite73 @shamil98 @shrutipande9 @sirm3d @timaashorty @TheKylieeM @thtshortchick @TTP @UCLAGRIND @undeadknight26 @vicleon6 @veritos @VictoriaLeone @V1r3z @wolf1728 @whambamsham @workhard.playhard @Xmoses1 @xoxo_jenni @young&reckless<3 @zeeporo @zaragoza96 any of you guys have an idea? :)

OpenStudy (anonymous):

Awww i'm sorryyy :( I'm stumped ! Good luck though :)

OpenStudy (anonymous):

Im horrible in math, sorry.

OpenStudy (nuke_ur_friend):

dont no sorry :( but good luck

jigglypuff314 (jigglypuff314):

umm... this is a lot of rules to explain :/ lets start with that 2/9 = 0.22222... 4/9 = 0.44444... 5/9 = 0.55555... see the pattern? so for 1/9 =

OpenStudy (anonymous):

0.1111111111 haha :)

jigglypuff314 (jigglypuff314):

yes :) and so 1/9 = 0.1 (with a line over the 1)

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

i dont know i srry :/

jigglypuff314 (jigglypuff314):

now for percents :P (number)% = (number)/100 so like 55% = 55/100 = 11/20 110% = 110/100 = (simplify) then it works the other way around as well... 4/25 = 16/100 = 0.16 = 16%

OpenStudy (anonymous):

you're good at this! haha

jigglypuff314 (jigglypuff314):

I'm not the only one good at this, there are other higher level people, who just don't want to spend their time explaining :P

OpenStudy (anonymous):

I don't know im srry

jigglypuff314 (jigglypuff314):

2/3 = 6/9 = remember the divided by 9 rules? :P

OpenStudy (anonymous):

2/25? is that wrong?

jigglypuff314 (jigglypuff314):

remember how 2/9 = 0.2222 4/9 = 0.4444 so what's 6/9 = ?

jigglypuff314 (jigglypuff314):

also, to convert decimals to percents (decimal) * 100 = ___% so like 0.375 to percent would be 0.375*100 = 37.5% then 0.125 to percent would be 0.125*100 = ____%

jigglypuff314 (jigglypuff314):

to convert decimals to fractions know that 0.1234 ^ tenths place ^ hundredths place ^ thousandths place so like 0.35 ^ hundredths place so 0.35 to fraction would be 35/100 which can be simplified so for 0.08 would be 8/100 which can be simplified and also for 0.125 ^ thousandths place so it would be 125/1000 which can be simplified :)

OpenStudy (zeeporo):

were you able to match everything?

OpenStudy (firejay5):

2/3 = 0.66 1/9 = 0.11 60% = 3/5 0.35 = 7/20 1/80 = 1.25% 4/25 = 16% 110% = 1.1 125% & 1 1/4 = 1.25 0.08 = 2/25 0.125 = 1/80

jigglypuff314 (jigglypuff314):

Firejay5 's answers aren't entirely correct... yes 1/80 = 1.25% but when converting 1.25% to decimal 1.25/100 = 0.0125 therefore his last match "0.125 = 1/80" was not correct 0.125 = 1/8

OpenStudy (firejay5):

1/80 in decimal form = 0.0125 and 1.25% in decimal form = 0.0125 not 1/8 recheck your calculation @jigglypuff314